# Multi-stage tomography based on eigenanalysis for high-dimensional dense unitary quantum processes

### v1

@@ Abstract
Quantum Process Tomography (QPT) methods aim at identifying, i.e. estimating, a quantum process. QPT is a major quantum information processing tool, since it especially allows one to experimentally characterize the actual behavior of quantum gates, that may be used as the building blocks of quantum computers. We here consider unitary, possibly dense (i.e. without sparsity constraints) processes, which corresponds to isolated systems. Moreover, we aim at developing QPT methods that are applicable to a significant number of qubits and hence to a high state space dimension, which allows one to tackle more complex problems. Using the unitarity of the process allows us to develop methods that first achieve part of QPT by performing an eigenanalysis of the estimated density matrix of a process output. Building upon this idea, we first develop a class of complete algorithms that are single-stage, i.e. that use only one eigendecomposition. We then extend them to multiple-stage algorithms (i.e. with several eigendecompositions), in order to address high-dimensional state spaces while being less limited by the estimation errors made when using an arbitrary given Quantum State Tomography (QST) algorithm as a building block of our overall methods. We first propose two-stage methods and we then extend them to dichotomic methods, whose number of stages increases with the considered state space dimension. The relevance of our methods is validated by means of simulations. Single-stage and two-stage methods first yield the following results. Just running them with standard PC and software already makes it possible to evaluate their performance for up to 13 qubits, i.e. with state space dimensions up to a few thousands. This shows their attractiveness in terms of accuracy and proves that they solve the core of the dense QPT problem in a very limited time frame. This also shows that moving from single-stage to two-stage methods allows one to decrease by a factor up to 36 the Normalized Mean Square Error (NMSE) obtained when estimating the process matrix. Moreover, for the lowest QST error magnitude considered in these tests, even for the highest analyzed state space dimensions, our two-stage methods yield an NMSE limited to about $10^{-2}$. Multi-stage dichotomic methods may then yield a higher accuracy (NMSE decreased by a factor up to 10 as compared with two-stage methods and NMSE limited to about $10^{-3}$ in almost the same conditions as above) but at the expense of a higher computational load.
@@ I Introduction
Quantum process tomography (QPT) is a major quantum data processing tool, since it especially allows one to characterize the actual behavior of quantum gates, that may be used as the building blocks of quantum computers. QPT was especially  [^1] See also [3] p. 398 for the other earliest references. introduced in 1997 in [2] and then extended e.g. in [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14].
QPT is a “‘data-driven” approach, and can thus be considered as a type of quantum machine learning (QML) problem (see [15] for a taxonomy of QML problems), because the input-output transform performed by the considered process is generally inferred from the values of input quantum states and/or from measurement results for output quantum states of that process. More precisely, usual QPT methods require one to fix a priori and hence know the above input “values” (i.e. states) of the process and to use information about the corresponding output “values” (i.e. states), that is derived from measurements performed at the output of the considered process. This corresponds to the supervised, i.e. non-blind, version of machine learning [16]. In contrast, the other extreme case is the one which requires the most limited knowledge about the process input whereas output measurements are still used, and this corresponds to the unsupervised, i.e. blind, QPT (BQPT) configuration [16]. BQPT methods were introduced in [17] and then especially developed in [18, 19, 16]. In that case, the individual input state values are completely unknown, that is, neither fixed a priori nor estimated by means of measurements: in [17, 18], the input consists of unknown unentangled deterministic-coefficients pure states (DCPS); in [19, 16], it consists of random-coefficient pure states (RCPS) and the information available about these RCPS is restricted to some of their global, i.e. statistical, properties. DCPS are the usual form of pure states of quantum physics, whereas RCPS were introduced in [20] and then discussed in various related papers, especially including [21].
@@
As detailed below, the different parts and variants of the QPT algorithms proposed in the present paper cover all above three types of learning. Some of them require their input states to be known, so that these states may be fixed a priori (supervised learning) or estimated from measurements (semi-supervised learning). Other parts and variants of our algorithms use input states whose values are not known, but that are required to meet a given condition (unsupervised learning). Besides, part of the proposed algorithms use “pure states” (in the usual sense, i.e. DCPS; this is not stated again in this paper), whereas other use statistical mixtures.
Within the above framework, we focus on an original case which combines the following two features. First, the considered process is constrained to be unitary, which corresponds to addressing a system that is isolated from its environment. Second, we are mainly interested in the case when the process has a high dimension, i.e. when it involves a significant number of qubits, which allows one to tackle more complex problems. We stress that we do not set any other conditions on the considered process. In particular, we allow (but do not request) the matrix that represents that process in the considered basis to be a so-called dense, i.e. non-sparse, matrix (we here use the standard terminology, i.e. a sparse matrix is a matrix that has a large percentage of zeros: see e.g. [23]). In contrast, some approaches from the literature for QPT and the related topic of Hamiltonian estimation request sparsity (see e.g. [24, 13]), but this was considered to be less relevant than unitarity in [22] and we explained above when the unitarity assumption is relevant.
@@
where $U$ is the unitary matrix that represents the process to be estimated in the considered basis, $^{{\dagger}}$ is the transconjugate (i.e. Hermitian transpose) and $\rho_{1}$ is the density matrix of the first input mixed state that yields the output state $\rho_{2}$. The idea is then that, when $\rho_{1}$ is selected to be diagonal, (1) is an eigendecomposition of $\rho_{2}$. Therefore, in practice, starting from an estimate of $\rho_{2}$ derived from measurements, its eigendecomposition may be hoped to yield a result directly related to an estimate of $U$. This idea will then have to be refined further in this paper because: 1) eigendecomposition yields indeterminacies (defined below) and 2) fluctuations in output density matrix estimation limit the performance of this preliminary approach and then lead us to extend it in this paper.
The above preliminary description shows that the type of approaches proposed in this paper has the following connections with respect to works that were previously reported both in the classical and quantum domains. QPT may be considered as the quantum version of so-called (classical) system identification [3]. The latter classical problem consists of estimating the parameter values of a transform from the input to the output of a system. That system may have multiple inputs, so that it then performs a combination of them. In the classical framework, such a combination is called a mixture. The classical problem then considered in the literature is therefore also referred to as mixture identification and it especially includes blind mixture identification, or BMI [25, 26, 27, 28]. This terminology should be used with care when also considering the quantum domain, because the above classical “mixtures” have nothing to do with the “statistical mixtures”, i.e. “mixed states”, faced in the quantum domain and used further in this paper. Only considering previous works in the classical domain at this stage of our discussion, the above-mentioned BMI has a close relationship with source separation, especially including blind source separation, or BSS [29, 30, 26, 27, 31, 32, 33]). BSS does not focus on identifying the transform performed by the above, i.e. direct, system itself, but mainly on deriving an estimate of the inverse of that transform, so as to apply it to the outputs of the above direct system and to thus restore the unknown source signals that form the input of the direct system. This relationship between BMI and BSS even appears in the name of the well-known SOBI method [34], which stands for “Second-Order Blind Identification”, whereas it is claimed to be a BSS method. The term “Second-Order” here refers to the fact that this method applies to random signals and uses the statistical properties of the measured signals, more precisely their usual second-order statistical parameters that consist of their covariance matrices involving different time delays between signals.
Moreover, the relationships between the above-mentioned SOBI method and related methods on the one hand and the present paper on the other hand not only concern the types of data processing problems they address, but also the methods that they use to solve these problems. More precisely, in classical BSS/BMI, when the transform performed by the direct system is linear (and memoryless) and its identification is performed by using the above-mentioned second-order statistics of the output signals of the direct system, an equation similar to (1) is obtained [35, 34, 28, 36, 37]. That classical counterpart of the structure of (1) was already exploited to perform classical BSS/BMI by means of several methods based on eigendecomposition. The real-valued version of the basic structure of (1) was first used in [35, 28] together with a required modified form of this equation (a related approach also appears in [37]). It was then extended in [34], by combining it with more than one modified form of this equation, to obtain a more robust method. Another version [36] aims at extending that type of approach from the standard case of stationary time-domain source signals to nonstationary time-domain sources. That version eventually uses complex-valued signals (whereas the mixing transform remains real-valued) because one moves to the Fourier domain.
@@
The first step of the investigation reported in the present paper (i.e. our single-stage methods) thus has some relationships with a few conventional classical signal processing methods. Besides, a few more recent papers, now dealing with quantum data, also consider (1) in the framework of QPT. However, they only have the following limited relationships with our investigation. Ref. [38] aims at deriving the number (and values) of quantum states that should be applied to the input of a quantum process to be able to identify that process. However, that paper does not focus on the same input values as in the present paper and, as also noted in [22], it does not provide an explicit QPT algorithm. Ref. [4] provides such algorithms, that are different from ours and that are tested only for a Hilbert space dimension $d=5$, whereas we consider dimensions up to $d=2^{13}\simeq 8000$.
@@
QPT algorithms may also be split in two classes, from the following point of view. Part of them first use the results of measurements at the output of the system by explicitly performing quantum state tomography (QST), i.e. by estimating the ket (for a pure state) or the density matrix (for a statistical mixture) that represents that ouput state. They then use these estimated output states (and the corresponding input states, their estimates or their properties) so as to infer the considered process. Other QPT methods instead directly use the results of the considered types of measurements. In the present paper, we consider the first class of QPT methods and we focus on the part dedicated to QPT itself, as opposed to QST, as detailed in Section V.
@@ II Proposed single-stage QPT methods
II.1 First part of the methods
@@
We here consider the situation when the first state applied to the analyzed process is a given statistical mixture, represented by the density matrix $\rho_{1}$. The process output is then also a mixed state and its density matrix $\rho_{2}$ may be expressed as (1). More specifically, we here introduce a first variant, in which the matrix $\rho_{1}$ is diagonal, whereas Appendix B addresses the more general case when this constraint is removed.
It should be noted that, when one considers that one is able to create a pure state, then, creating the type of state considered here is not a problem, as will now be shown. We here consider the approach initially used by von Neumann in [42] to define a statistical mixture: such a state is expressed as being composed of a set of pure states, each of them being associated with a probability of occurrence. Then, to create experimental data corresponding to the state defined by the diagonal density matrix $\rho_{1}$, one just has to repeatedly, randomly, select one of the pure states that form the considered basis of the state space, moreover selecting the $k$th basis state with a probability that is equal to the $k$th value on the diagonal of $\rho_{1}$. This simplicity of implementation is worth noting because, in constrast, the authors of [38, 4] consider that using a mixed input state is too complex (maybe because they have non-diagonal density matrices in mind), so that they choose to restrict themselves to pure states, which has the drawback of requesting many such multiqubit input states, as opposed to only two multiqubit input states in the complete approach that we here start to describe (namely the mixed state with a density matrix $\rho_{1}$ considered here, and then the pure state $\lvert\Psi_{1}\rangle$ in Section II.2). Beyond their relationship based on considering (1), our approach and the one in [38] are therefore quite different.
@@
The matrix $\rho_{2}$ is known to be Hermitian, since it is a density matrix (see e.g. [39] p. 283). Therefore, it is normal (see [43] p. 100;  [^2] Throughout this paper, the reader may refer to [65], instead of [43].) and hence unitarily diagonalizable: see Theorem 1 and comments in its proof (all theorems cited in the core of this paper are provided in Appendix A). The first part of the QPT method proposed below is based on the eigendecomposition of $\rho_{2}$ (because (1) is such an eigendecomposition, as detailed below). More precisely, in practice, what is initially available for performing an eigendecomposition is not the theoretical matrix $\rho_{2}$ but an estimate $\widehat{\rho}_{2}$ of $\rho_{2}$, provided by a QST algorithm. Therefore, if this estimate $\widehat{\rho}_{2}$ is not guaranteed to be Hermitian due to estimation errors, a first recommended preprocessing step of our method consists of deriving a Hermitian estimate $\widehat{\rho}_{3}$ from $\widehat{\rho}_{2}$. This may be achieved by computing the Hermitian part of $\widehat{\rho}_{2}$, that is (see [43] p. 109 or [31] p. 343 for it real-valued counterpart):
@@
After these preprocessing steps, the next step of our algorithm consists of performing an eigendecomposition of the Hermitian matrix $\widehat{\rho}_{4}$. This eigendecomposition problem has several solutions, corresponding to the indeterminacies of eigendecomposition detailed below. Any of these solutions is defined by two matrices, hereafter denoted as $D_{4}$ and $V_{4}$. $D_{4}$ is a diagonal matrix that contains the eigenvalues of $\widehat{\rho}_{4}$ in a arbitrary order: see e.g. [43] p. 102, that states “the diagonal entries of any diagonalization of [a matrix] are the eigenvalues $\lambda_{1},\dots,\lambda_{n}$ (in some order)”; about the ordering of eigenvalues in Schur’s decomposition and the case of a normal matrix, see also [46] p. 313, 314 and 365 (in practice, for instance, the Matlab software that we used for eigendecomposition does provide eigenvalues in an arbitrary order and this is stated in its documentation). Moreover, each column of $V_{4}$ is an eigenvector of $\widehat{\rho}_{4}$ (see (4) and Theorem 2) and $V_{4}$ is here unitary (because $\widehat{\rho}_{4}$ is unitarily diagonalizable), hence $V_{4}^{-1}=V_{4}^{{\dagger}}$. These results $V_{4}$ and $D_{4}$ of the eigendecomposition are linked to the initial matrix $\widehat{\rho}_{4}$ as follows:
@@
In our first type of QPT methods described here, we constrain all diagonal values of the user-defined matrix $\rho_{1}$ to be different. This means that all eigenvalues in the eigendecompositions of $\rho_{2}$ and hence $\widehat{\rho}_{4}$ (ignoring estimation errors) are different. Therefore, each of them is associated with a different one-dimensional eigenspace, i.e. all eigenvectors corresponding to a given eigenvalue are along the same direction and the directions associated with two different eigenvalues are different (this is true for any complex square matrix: see [43] p. 47) and even orthogonal: see Theorem 3. Each eigenvalue is thus coupled with a single eigenvector in any considered eigendecomposition. The first indeterminacy of the eigendecomposition of $\widehat{\rho}_{4}$ is then a possible coupled permutation of 1) the columns of the matrix $V_{4}$ obtained as a first quantity towards the estimation of $U$ in this eigendecomposition and 2) the values along the diagonal of the matrix $D_{4}$, due to the above-mentioned arbitrary order of the latter diagonal values. This permutation problem may be solved as follows. As stated above, we constrain all diagonal values of $\rho_{1}$ to be different (at this stage of the discussion, we restrict ourselves to a theoretical approach, where we only need the eigenvalues to be different; further in this paper, we will discuss the gap required between them, when considering practical implementations of our approach). Moreover, we constrain all diagonal values of $\rho_{1}$ to be placed on that diagonal according to a given order. Hereafter, we will consider the case when they are in decreasing order. With respect to the above-defined machine learning terminology, this part of our algorithm is thus unsupervised: it does not require the exact value of the input state, i.e. the input density matrix $\rho_{1}$, to be known but it sets limited conditions on some of its properties.
@@
Thanks to the above approach, the only indeterminacy that remains in $\widehat{U}_{2}$ with respect to $U$ is an unknown complex scale factor in each of its columns, since any such column belongs to the adequate one-dimensional eigenspace and is therefore defined up to a complex scale factor. Moreover, we create $\widehat{U}_{2}$ so that each of its columns has unit norm, i.e., if the considered eigendecomposition algorithm does not directly guarantee that this condition is met, then we postprocess the value of $\widehat{U}_{2}$ that it yields, by dividing each column of $\widehat{U}_{2}$ by its norm. After this postprocessing, the scale factor of each column of $\widehat{U}_{2}$, with respect to the column of $U$ that has the same index, reduces to a unit-modulus factor, i.e. a phase factor, because $U$ is unitary and therefore all its columns have unit norm, as those of $\widehat{U}_{2}$ (this phase indeterminacy of eigendecomposition is e.g. mentioned in the online documentation of the Matlab software that we used). $\widehat{U}_{2}$ then reads
@@
The second part of our algorithm, presented below in Section II.2, aims at determining all phase factors $e^{i\phi_{1}},\dots,e^{i\phi_{d}}$, up to a global phase factor. It should be noted that, whatever the considered QPT method, $U$ can only be defined up to a single, i.e. global, phase factor, since this phase factor then appears in the resulting expression of the ket associated with a pure output state of the considered process, and indeed the ket for a pure quantum state can be meaningfully defined only up to a global phase factor. Our methods can therefore only aim at estimating $U$ up to a single, i.e. global, phase factor. In contrast, having a set of uncontrolled phase factors, with one factor per column of the estimate of $U$ (as in $\widehat{U}_{2}$) is not acceptable, because two such estimates with arbitrary phase factors do not represent the same behavior: the resulting output kets are completely different.
In the first part of our method defined in the present section, we preferably use the value of $\rho_{1}$ defined as follows. The values on its diagonal are the eigenvalues of the considered density operator $\rho_{2}$, they are real and non-negative and their sum is equal to one (see e.g. [3] p. 101). As shown above, the proposed method assumes that these eigenvalues are different and ordered, and it is based on reordering the eigenvalues in the eigendecomposition of the estimated output density matrix and on permuting the eigenvectors accordingly. Therefore, if using quite close actual eigenvalues when choosing $\rho_{1}$, one may fear that the corresponding estimated eigenvalues obtained for the output density matrix be permuted due to estimation errors, so that the corresponding eigenvectors may then be permuted accordingly. This is a major problem, because the intermediate estimate $\widehat{U}_{2}$ related to the considered process, obtained at the output of this first step of our QPT method, is primarily based on these eigenvectors. In this paper, we handle this problem as follows: first, for the method imposed in the present section, we select the value of $\rho_{1}$ in order to limit that problem to the greatest extent that is possible within the frame of that method; then, in Sections III and IV, we propose more advanced methods that allow one to further reduce that problem. So, for the method considered here, we choose eigenvalues that are uniformly distributed (i.e. multiples of the same step, with zero excluded), so as to maximize the gap between any two adjacent values. Moreover constraining them to be positive, with a sum equal to one and in decreasing order as explained above, it is easily shown that one obtains the following values for the diagonal elements $\rho_{1,kk}$ of $\rho_{1}$:
II.2 Second part of the methods
Thanks to the limited number of unknowns (namely the phase factors $e^{i\phi_{1}},\dots,e^{i\phi_{d}}$) that remain in the matrix $\widehat{U}_{2}$ of (6) obtained in Section II.1, a single and pure state is then sufficient for estimating these unknowns. This yields the first variant of the second part of our methods, presented here. However, one may instead use a mixed state for estimating these unknowns, as shown in Appendix C.
We therefore here apply a known pure state, defined by a ket $\lvert\Psi_{1}\rangle$, at the input of the process (this part of the algorithm is therefore supervised, or semi-supervised if $\lvert\Psi_{1}\rangle$ is initially unknown but estimated by performing measurements for part of the available copies of that state). The corresponding pure state at the output of that process is defined by the ket

### v2

@@ Abstract
Quantum Process Tomography (QPT) methods aim at identifying, i.e. estimating, a quantum process. QPT is a major quantum information processing tool, since it especially allows one to experimentally characterize the actual behavior of quantum gates, that may be used as the building blocks of quantum computers. We here consider unitary, possibly dense (i.e. without sparsity constraints) processes, which corresponds to isolated systems. Moreover, we develop QPT methods that are applicable to a significant number of qubits and hence to a high state space dimension, which allows one to tackle more complex problems. Using the unitarity of the process allows us to develop methods that first achieve part of QPT by performing an eigenanalysis of the estimated density matrix of a process output. Building upon this idea, we first develop a class of complete algorithms that are single-stage, in the sense that they use only one eigendecomposition. We then extend them to multiple-stage algorithms (i.e. with several eigendecompositions), in order to address high-dimensional state spaces while being less limited by the estimation errors made when using an arbitrary given Quantum State Tomography (QST) algorithm as a building block of our overall methods. We first propose two-stage methods and we then extend them to dichotomic methods, whose number of stages increases with the considered state space dimension. The relevance of our methods is validated with simulations. Single-stage and two-stage methods efficiently apply up to 13 qubits on a standard PC (with 16 GB of RAM). Multi-stage methods yield an even higher accuracy (Normalized Mean Square Error down to $10^{-3}$).
Index Terms:
eigenanalysis of density matrix, high-dimensional quantum state space, quantum computing, quantum gate characterization, quantum machine learning (supervised / semi-supervised / unsupervised), quantum process tomography, quantum state tomography, unitary quantum gate
[…]
=-15pt
@@ I Introduction
Quantum process tomography (QPT) may be considered as the quantum counterpart of classical system identification [1], that is, system parameter estimation. It was mainly [^1] See also [1] p. 398 for the other earliest references. introduced in 1997 in [2] and then extended (together with quantum process learning) e.g. in [1, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27]. QPT is a major quantum data processing tool, since it especially allows one to characterize the actual behavior of quantum gates, that may be used as the building blocks of quantum computers. This gate-based (or circuit-based) approach builds upon the unitary transforms performed by quantum gates (see Chapter 4 of [1]). It is currently one of the most prominent quantum computing paradigms and various commercial products based on this paradigm are available. This e.g. includes IBM’s Qiskit environment (see [28, 29, 30] and especially [31] for the use of reversible unitary gates), Amazon’s Braket tool (see [32] especially for reversible transformations), and Mathworks’s software package (see [33] e.g. for their statement “Quantum gates represent reversible operations that transform the quantum state according to unitary matrices”). Due to the importance of the above-mentioned unitary processes, various methods dedicated to such processes have been proposed, for performing unitary quantum process estimation/learning (see e.g. [3, 22, 13, 23, 24, 25, 26, 27]) and e.g. for the related topic of the inversion and transposition of unknown unitary operations [34].
QPT is a “‘data-driven” approach, and can thus be considered as a type of quantum machine learning (QML) problem (see [35] for a taxonomy of QML problems), because the input-output transform performed by the considered process is generally inferred from the values of input quantum states and/or from measurement results for output quantum states of that process. More precisely, usual QPT methods require one to fix a priori and hence know the above input “values” (i.e. states) of the process and to use information about the corresponding output “values” (i.e. states), that is derived from measurements performed at the output of the considered process. This corresponds to the supervised version of machine learning, also called the non-blind version of machine learning [36]. In contrast, the other extreme case is the one which requires the most limited knowledge about the process input whereas output measurements are still used, and this corresponds to the unsupervised, or blind, QPT (BQPT) configuration [36]. BQPT methods were introduced in [37] and then especially developed in [38, 39, 36]. In that case, the individual input state values are completely unknown, that is, neither fixed a priori nor estimated by means of measurements: in [37, 38], the input consists of unknown unentangled deterministic-coefficients pure states (DCPS); in [39, 36], it consists of random-coefficient pure states (RCPS) and the information available about these RCPS is restricted to some of their global, i.e. statistical, properties. DCPS are the usual form of pure states of quantum physics, whereas RCPS were introduced in [40] and then discussed in various related papers, especially including [41].
@@
As detailed below, the different parts and variants of the QPT algorithms proposed in the present paper cover all above three types of learning. Some of them require their input states to be known, so that these states may be fixed a priori (supervised learning) or estimated from measurements (semi-supervised learning). Other parts and variants of our algorithms use input states whose values are not known, but that are required to meet a given condition (unsupervised learning). Besides, part of the proposed algorithms use “pure states”, where we employ the term “pure states” in the above-defined usual sense, that is, DCPS. In contrast, other proposed algorithms use statistical mixtures.
Within the above framework, we focus on an original case which combines the following two features. First, the considered process is constrained to be unitary, which corresponds to addressing a system that is isolated from its environment. Second, we are mainly interested in the case when the process has a high dimension, because it involves a significant number of qubits. This allows one to tackle more complex problems. We stress that we do not set any other conditions on the considered process. In particular, we allow (but do not request) the matrix that represents that process in the considered basis to be a so-called dense, i.e. non-sparse, matrix (we here use the standard terminology: a sparse matrix is a matrix that has a large percentage of zeros; see e.g. [42]). In contrast, some approaches from the literature for QPT and the related topic of Hamiltonian estimation request sparsity (see e.g. [43, 12]), but this was considered to be less relevant than unitarity in [22] and we explained above when the unitarity assumption is relevant.
@@
where $U$ is the unitary matrix that represents the process to be estimated in the considered basis, $^{{\dagger}}$ is the transconjugate (i.e. Hermitian transpose) and $\rho_{1}$ is the density matrix of the first input mixed state that yields the output state $\rho_{2}$. The idea is then that, when the density matrix $\rho_{1}$ is selected to be diagonal in the considered and fixed basis (computational basis), Eq. (1) is an eigendecomposition of $\rho_{2}$. Therefore, in practice, starting from an estimate of $\rho_{2}$ derived from measurements, its eigendecomposition may be hoped to yield a result directly related to an estimate of $U$. This idea will then have to be refined further in this paper because: 1) eigendecomposition yields indeterminacies (defined below) and 2) fluctuations in output density matrix estimation limit the performance of this preliminary approach and then lead us to extend it in this paper.
@@
The above preliminary description shows that the type of approaches proposed in this paper has some connections with works that were previously reported both in the classical and quantum domains. The connections with classical approaches are discussed in Appendix A, whereas those with the quantum domain may be defined as follows. A few recent papers dealing with quantum data also consider (1) in the framework of QPT. However, they only have the following limited relationships with our investigation. Ref. [44] aims at deriving the number (and values) of quantum states that should be applied to the input of a quantum process to be able to identify that process. However, that paper does not focus on the same input values as in the present paper and, as also noted in [22], it does not provide an explicit QPT algorithm. Ref. [3] provides such algorithms, that are different from ours and that are tested only for a Hilbert space dimension $d=5$, whereas we consider dimensions up to $d=2^{13}\simeq 8000$.
@@
QPT algorithms may also be split in two classes, from the following point of view. Part of them first use the results of measurements at the output of the system by explicitly performing quantum state tomography (QST). This means that they estimate the ket (for a pure state) or the density matrix (for a statistical mixture) that represents that ouput state. They then use these estimated output states (and the corresponding input states, their estimates or their properties) so as to infer the considered process. Other QPT methods instead directly use the results of the considered types of measurements. In the present paper, we consider the first class of QPT methods and we focus on the part dedicated to QPT itself, as opposed to QST, as detailed in Section V.
@@ II Proposed single-stage QPT methods
II-A First part of the methods
@@
We here consider the situation when the first state applied to the analyzed process is a given statistical mixture, represented by the density matrix $\rho_{1}$. The process output is then also a mixed state and its density matrix $\rho_{2}$ may be expressed as (1). More specifically, we here introduce a first variant, in which the matrix $\rho_{1}$ is diagonal (see Appendix B about how to create it), whereas Appendix G addresses the more general case when this constraint is removed.
@@
The matrix $\rho_{2}$ is known to be Hermitian, since it is a density matrix (see e.g. [45] p. 283). Therefore, it is normal (see [48] p. 100;  [^2] Throughout this paper, the reader may refer to [49], instead of [48].) and hence unitarily diagonalizable: see Theorem 1 and comments in its proof (all theorems cited in the core of this paper are provided in Appendix F). The first part of the QPT method proposed below is based on the eigendecomposition of $\rho_{2}$ (because (1) is such an eigendecomposition, as detailed below). More precisely, in practice, what is initially available for performing an eigendecomposition is not the theoretical matrix $\rho_{2}$ but an estimate $\widehat{\rho}_{2}$ of $\rho_{2}$, provided by a QST algorithm (we here use standard notations, e.g. as in [50, 51, 52]: an estimate of an unknown, deterministic, scalar, vector or matrix $\theta$ is denoted as $\widehat{\theta}$). Therefore, our QPT method preferably starts with two preprocessing steps, that consist of first deriving a Hermitian estimate $\widehat{\rho}_{3}$ from $\widehat{\rho}_{2}$ and then a unit-trace estimate $\widehat{\rho}_{4}$ from $\widehat{\rho}_{3}$ (see details in Appendix C), unless the considered QST algorithm already guarantees that $\widehat{\rho}_{2}$ meets these two properties.
@@
Then, the core of our QPT algorithm performs an eigendecomposition of the Hermitian matrix $\widehat{\rho}_{4}$. This eigendecomposition problem has several solutions, corresponding to the indeterminacies of eigendecomposition detailed below. Any of these solutions is defined by two matrices, hereafter denoted as $D_{4}$ and $V_{4}$. $D_{4}$ is a diagonal matrix that contains the eigenvalues of $\widehat{\rho}_{4}$ in a arbitrary order (see e.g. [48] p. 102). Moreover, each column of $V_{4}$ is an eigenvector of $\widehat{\rho}_{4}$ (see (2) and Theorem 2) and $V_{4}$ is here unitary (because $\widehat{\rho}_{4}$ is unitarily diagonalizable), hence $V_{4}^{-1}=V_{4}^{{\dagger}}$. These results $V_{4}$ and $D_{4}$ of the eigendecomposition are linked to the initial matrix $\widehat{\rho}_{4}$ as follows:
@@
In our first type of QPT methods described here, we constrain all diagonal values of the user-defined matrix $\rho_{1}$ to be different. This means that all eigenvalues in the eigendecompositions of $\rho_{2}$ and hence $\widehat{\rho}_{4}$ (ignoring estimation errors) are different. Therefore, each of them is associated with a different one-dimensional eigenspace. In other words, all eigenvectors corresponding to a given eigenvalue are along the same direction and the directions associated with two different eigenvalues are different (this is true for any complex square matrix: see [48] p. 47) and even orthogonal: see Theorem 3. Each eigenvalue is thus coupled with a single eigenvector in any considered eigendecomposition. The first indeterminacy of the eigendecomposition of $\widehat{\rho}_{4}$ is then a possible coupled permutation of 1) the columns of the matrix $V_{4}$ obtained as a first quantity towards the estimation of $U$ in this eigendecomposition and 2) the values along the diagonal of the matrix $D_{4}$, due to the above-mentioned arbitrary order of the latter diagonal values. This permutation problem may be solved as follows. As stated above, we constrain all diagonal values of $\rho_{1}$ to be different (at this stage of the discussion, we restrict ourselves to a theoretical approach, where we only need the eigenvalues to be different; further in this paper, we will discuss the gap required between them, when considering practical implementations of our approach: see Appendix D for the method considered in the present section). Moreover, we constrain all diagonal values of $\rho_{1}$ to be placed on that diagonal according to a given order. Hereafter, we will consider the case when they are in decreasing order. With respect to the above-defined machine learning terminology, this part of our algorithm is thus unsupervised: it does not require the exact value of the input state, i.e. the input density matrix $\rho_{1}$, to be known but it sets limited conditions on some of its properties.
@@
Thanks to the above approach, the only indeterminacy that remains in $\widehat{U}_{2}$ with respect to $U$ is an unknown complex scale factor in each of its columns, since any such column belongs to the adequate one-dimensional eigenspace and is therefore defined up to a complex scale factor. Moreover, we create $\widehat{U}_{2}$ so that each of its columns has unit norm. To this end, if the considered eigendecomposition algorithm does not directly guarantee that this unit-norm condition is met, we postprocess the value of $\widehat{U}_{2}$ that it yields, by dividing each column of $\widehat{U}_{2}$ by its norm. After this postprocessing, the scale factor of each column of $\widehat{U}_{2}$, with respect to the column of $U$ that has the same index, reduces to a unit-modulus factor, i.e. a phase factor, because $U$ is unitary and therefore all its columns have unit norm, as those of $\widehat{U}_{2}$. $\widehat{U}_{2}$ then reads
@@
The second part of our algorithm, presented below in Section II-B, aims at determining all phase factors $e^{i\phi_{1}},\dots,e^{i\phi_{d}}$, up to a global phase factor (a process matrix $U$ can only be defined up to a single, i.e. global, phase factor).
II-B Second part of the methods
Thanks to the limited number of unknowns (namely the phase factors $e^{i\phi_{1}},\dots,e^{i\phi_{d}}$) that remain in the matrix $\widehat{U}_{2}$ of (4) obtained in Section II-A, a single and pure state is then sufficient for estimating these unknowns. This yields the first variant of the second part of our methods, presented here. However, one may instead use a mixed state for estimating these unknowns, as shown in Appendix H.
We therefore here apply a known pure state, defined by a ket $\lvert\Psi_{1}\rangle$, at the input of the process (this part of the algorithm is therefore supervised, or semi-supervised if $\lvert\Psi_{1}\rangle$ is initially unknown but estimated by performing measurements for part of the available copies of that state; for the supervised case, Appendix E explains how $\lvert\Psi_{1}\rangle$ is chosen). The corresponding pure state at the output of that process is defined by the ket